OpenStudy (anonymous):

zeno takes 3hrs oonger to paint a floor than it takes lia when they work together it takes them 2 hrs how long would each take to do the job alone?

OpenStudy (anonymous):

1 room in 6 hours and 1 room in 3 hours or 1/3 and 1/6, 1/3 1/6+1/3 = 3/6 or 1/2, 1 room in 2 hours

OpenStudy (anonymous):

so, lia takes 3 hours and zeno takes 6 hours

OpenStudy (anonymous):

Better yet, lia can do 1/3 of the room in an hour and zeno can do 1/6 of the room in an hour, so in 2 hours lia can do 2/3 of the room while zeno does 2/6=1/3 of the room in the same 2 hours 2/3 + 1/3 = 1 complete room